OpenStudy (anonymous):

CAN SOMEONE PLEASE HELP ME FIND THE ABSOLUTE MINIMUM AND ABSOLUTE MAXIMUM OF THIS EQUATION? IM SO CONFUSED

OpenStudy (anonymous):

\[y=x \sqrt{25-x^2}\]

OpenStudy (anonymous):

\[Let s=y ^{2}=x ^{2}\left( 25-x ^{2} \right)=25x ^{2}-x ^{4}\] \[\frac{ ds }{ dx }=50x-4x ^{3}\]

OpenStudy (anonymous):

so what do i do next?

OpenStudy (anonymous):

\[\frac{ d ^{2}s }{dx ^{2} }=50-12x ^{2}\]

OpenStudy (anonymous):

\[find stationary values by plugging \frac{ ds }{dx }=0,then for these values find y \] you can check for maxima and minima by plugging in \[\frac{ d ^{2}s }{dx ^{2} },if \frac{ d ^{2} s}{dx ^{2} }>0,then relative minima,\]

OpenStudy (anonymous):

\[if \frac{ d ^{2}s }{ dx ^{2} }<0,then relative maxima.\]

OpenStudy (anonymous):

i have no idea what any of that means. i've never learned any of that.. we're supposed to do it on the graphing calculator

OpenStudy (anonymous):

y exists only if \[25-x ^{2}\ge 0,or 25\ge x ^{2},x ^{2}\le 25,\left| x \right|\le 5,-5 \le x \le 5\]

OpenStudy (anonymous):

\[50x-4x ^{3}=0,2x \left( 25-2x ^{2} \right)=0,\]

OpenStudy (anonymous):

Either 2x=0,gives x=0 \[or 25-2x ^{2}=0,x=\pm \frac{ 5 }{\sqrt{2} }\]

OpenStudy (anonymous):

\[so x=-5,\frac{ -5 }{\sqrt{2} },0,\frac{ 5 }{\sqrt{2} },5\]

OpenStudy (anonymous):

find y for these values and find absolute maximum and absolute minimum value.

OpenStudy (anonymous):

how would i find y

OpenStudy (anonymous):

\[put the values of x \in y=x \sqrt{25-x ^{2}}, one by one.\]

OpenStudy (anonymous):

show me your work after solving for x=-5

OpenStudy (anonymous):

so i would make x=0? y=0\[\sqrt{25-0^2}\]

OpenStudy (anonymous):

\[when x=-5,y=-5\sqrt{25-25}=-5*0=0\]

OpenStudy (anonymous):

put other 4 values

OpenStudy (anonymous):

\[when x=\frac{ -5 }{\sqrt{2}} ,y=-\frac{ -5 }{\sqrt{2} }\sqrt{25-\frac{ 25 }{2 }}=\frac{ -5 }{ \sqrt{2} }*\frac{ 5 }{\sqrt{2} }=\frac{ -25 } { 2 }\]

OpenStudy (anonymous):

similarly find y for x=0,x=5/sqrt2,x=5